# Quasi-steady emission from repeating fast radio bursts can be explained by magnetar wind nebula

### v1

@@ 5.3 Rotation and magnetic powered models
Figure 7: Effect of NS age ($t_{\rm age}$) on the radio SED of persistent emission from FRB 121102 (190520) [201124] source is shown in the left (center) [right] panel. The fluxes for synchrotron emission are shown for magnetar-flare-powered energy injection model with $\epsilon_{B}=0.1$, $\gamma_{b}=10^{3}$, $q_{1}=q_{2}=2.0$. For each source, we fix $P_{i}=300\,{\rm ms}$ and $B_{\rm dip}\sim(0.01-0.1)\,B_{\rm int}\approx 5\times 10^{14}-5\times 10^{15}\,{\rm G}$ to vary the NS age within a range that includes the best-fit $t_{\rm age}$. In each panel, the SED data from radio observations is shown with filled circles and the results for USSN/CCSN progenitors are shown using solid/dashed curves.
@@
Figure 8: Radio light curves for persistent emission associated with FRB 121102 (190520) [201124] are shown in the left (center) [right] panel. Corresponding data at various frequencies is shown using unfilled circles for each source. The solid and dashed curves here denote the results for USSN and CCSN progenitors, respectively. We fix the NS parameters ($B_{\rm dip}$, $P_{i}$, $t_{\rm age}$) to their best-fit values as obtained from Figure 7 for the magnetar-flare-powered model.
@@
  FRB 121102 USSN ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) $25\,{\rm yr}$ $3\times 10^{15}\,{\rm G}$
@@
FRB 190520 USSN ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) $40\,{\rm yr}$ $5\times 10^{15}\,{\rm G}$
@@
FRB 201124 USSN ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) $12.5\,{\rm yr}$ $5\times 10^{14}\,{\rm G}$
@@
Energy injected by the rotating magnetar drives the MWN and SN ejecta to evolve together, and the observed quasi-steady emission is powered by the synchrotron radiation generated from the magnetized nebula. For young rapidly rotating NS, the rotational energy is the primary energy reservoir that powers the magnetised nebula and it has been studied in detail for Galactic PWNe (Tanaka & Takahara, 2010, 2013). For a typical NS with mass $M_{\rm ns}=1.4\,M_{\odot}$ and radius $R_{\rm ns}=12\,{\rm km}$, the rotational energy is given by $\mathcal{E}_{\rm rot,i}\approx(1.9\times 10^{52}\,{\rm erg})\,P_{i,-3}^{-2}$. However, for a magnetar of age $t_{\rm age}\gg t_{\rm sd}\approx(0.12\,{\rm yr})\,B_{\rm dip,13}^{-2}P_{\rm i,-3}^{2}$, the central NS may spin down considerably over time such that its interior magnetic energy, $\mathcal{E}_{\rm B,int}\approx B_{\rm int}^{2}R_{\rm ns}^{3}/6=(3\times 10^{49}\,{\rm erg})\,B_{\rm int,16}^{2}$, is comparable or even larger than $\mathcal{E}_{\rm rot,i}$.
@@
In our analysis, we incorporate models for rotational energy injection proposed by Kashiyama & Murase (2017) and energy injection due to magnetar flares based on Margalit & Metzger (2018), to compute the total energy injected into the magnetar wind nebula (see equation 2.2.1). However, the microphysical parameters $\epsilon_{B}$, $\gamma_{b}$, $q_{1}$ and $q_{2}$ that determine the actual energy injection rate into the radiating electrons are given by equation (4). In the previous sections, we considered a rotation-powered model with $\epsilon_{B}=0.01$, $\gamma_{b}=10^{5}$, $q_{1}=1.5$ and $q_{2}=2.5$. The corresponding best-fit NS parameters ($t_{\rm age}$, $B_{\rm dip}$, $P_{i}$) are listed in Table 1 for USSN and CCSN progenitor models, based on the radio SEDs obtained for all three sources in Figures 3 and 4.
@@
We now consider the scenario where the energy injection from magnetar flares exceeds that from NS rotation i.e. $\mathcal{E}_{\rm B,int}\gtrsim\mathcal{E}_{\rm rot,i}$, with $P_{i}=300\,{\rm ms}$, $B_{\rm int}=5\times 10^{16}\,{\rm G}$ and $B_{\rm dip}\sim 5\times 10^{14}-5\times 10^{15}\,{\rm G}$. We study the effect of varying $t_{\rm age}$ on the radio SED of synchrotron emission for the magnetar-flare-powered model with $\epsilon_{B}=0.1$, $\gamma_{b}=10^{3}$ and $q_{1}=q_{2}=2.0$. The results for R1/R2/R3 are shown in the left/center/right panel of Figure 7, with solid/dashed curves denoting USSN/CCSN progenitors and filled circles corresponding to the radio data for each source. As for the rotation-powered model in Figure 3, the late-time ($t_{\rm age}\gg t_{\rm sd}$) and low-energy ($\nu\lesssim 1\,{\rm GHz}$) synchrotron flux is suppressed due to significant adiabatic losses and SSA, respectively. However, the high-energy radio flux typically attains smaller values compared to the rotation-powered model to then gradually fall off. This is expected as the energy injection into particles occurs at smaller $\gamma_{b}\sim 5\times 10^{3}$ for the magnetar-flare-powered model. We find that the USSN progenitor model better explains the low-energy ($\nu\lesssim 1\,{\rm GHz}$) radio SED for both R1 and R2, corresponding to best-fit $t_{\rm age}\approx 25\,{\rm yr}$ and $\approx 40\,{\rm yr}$, respectively. On the contrary, the radio SED for R3 is only explained by the CCSN model with a best-fit $t_{\rm age}\approx 12.5\,{\rm yr}$ while accounting for the large suppression in radio flux at smaller energies.
@@
For the magnetar-flare-powered model, we show the radio light curves for the persistent emission associated with each FRB source in Figure 8. We use the best-fit $B_{\rm dip}$ and $P_{i}$ values obtained from radio SED fits in Figure 7 for each FRB source, with the corresponding NS age set to $t_{\rm age}=25\,(40)\,[12.5]\,{\rm yr}$. In contrast to the results obtained in Figure 5 for the rotation-powered model, we find that the flux predictions from the USSN model are favored by the radio data, especially at smaller frequencies, for both R1 and R2. In case of R3, the radio observations at smaller frequency $\nu\approx 6\,{\rm GHz}$ favor the CCSN model over the USSN progenitor model. These results are consistent with those obtained from the radio SED fits for the magnetar-flare-powered model ($\epsilon_{B}=0.1$, $\gamma_{b}=10^{3}$, $q_{1}=q_{2}=2.0$) in Figure 7. The best-fit NS parameters ($t_{\rm age}$, $B_{\rm dip}$) for the magnetar-flare-powered scenario with $P_{i}\approx 300\,{\rm ms}$ are listed in Table 2, for both USSN and CCSN models, based on the radio SED and light curve results shown in Figure 7 and 8, respectively.
6 Summary & Discussions
@@
Magnetars and rapidly rotating pulsars and magnetars have been studied extensively as central engines of SLSNe and FRBs, as they can naturally explain the quasi-steady synchrotron emission from their nascent pulsar/magnetar wind nebulae. These central engines can provide a unified picture of SLSNe, stripped-envelope SNe, FRBs as well as long GRBs. Non-thermal nebular emission has been observed for many Galactic PWNe, which suggests that a significant fraction of the wind magnetic energy can be utilized for accelerating particles. Murase et al. (2016) proposed quasi-steady synchrotron emission as a potential probe of FRB progenitors and also explored their possible connection to pulsar/magnetar-driven SNe including SLSNe.
@@
Till date, three repeating FRBs namely FRB 121102, FRB 190520 and FRB 201124 have been localized and associated with compact persistent radio sources (Chatterjee et al., 2017; Niu et al., 2022; Bruni et al., 2023). In this study, we consider the scenario whereby quasi-steady emission from such FRB sources are powered by young rapidly rotating magnetars that are embedded in a composite of magnetized wind nebula and SN ejecta. We have presented a detailed theoretical model here to numerically compute the radio SEDs for nebular emission by solving the time-dependent kinetic equations for photons and electrons in the MWN. To test the SNe progenitor models, we considered two cases: (a) an ultra-stripped SNe with $M_{\rm ej}=0.1\,M_{\odot}$ and $E_{\rm sn}=10^{50}\,{\rm erg}$, and (b) a conventional core-collapse SNe with $M_{\rm ej}=3.0\,M_{\odot}$ and $E_{\rm sn}=10^{51}\,{\rm erg}$.
@@
While rotational energy from a young NS is likely to be the primary reservoir that powers magnetized nebula, energy released from the dissipation of interior magnetic fields can be significant at later times $t_{\rm age}\gg t_{\rm sd}$ once the NS spins down. We estimated the total energy injected into the MWN based on the NS rotational energy injection model proposed by Kashiyama & Murase (2017). For the magnetar flare energy injection model, we adopted $B_{\rm int}=10^{16}\,{\rm G}$ and $t_{\rm inj}=0.6\,{\rm yr}$ as suggested by Margalit & Metzger (2018). In addition to the SNe progenitor and energy injection models, synchrotron flux from the magnetized nebula is determined by NS parameters such as $B_{\rm dip}\sim 10^{12}-10^{15}\,{\rm G}$, $P_{i}\sim 1-30\,{\rm ms}$ and $t_{\rm age}\gtrsim t_{\rm obs}$.
@@
We first estimated the allowed combinations of these parameters using physical constraints from: (a) energy injected into the magnetized nebula ($\varepsilon_{\rm nb,min}<\varepsilon_{\rm inj,max}$), (b) near-source DM contribution from the SN ejecta and MWN (${\rm DM}_{\rm ej+nb}<{\rm DM}_{\rm host}$), (c) non-attenuation of radio signal due to free-free absorption and synchrotron self-absorption ($\tau_{\rm ff},\,\tau_{\rm sa}\lesssim 1$), and (d) size of the magnetized nebula ($R_{\rm nb}\lesssim 10\,{\rm pc}$). The corresponding results for $B_{\rm dip}$, $P_{i}$ and $t_{\rm age}$ are shown in Figure 2. We showed that $\varepsilon_{\rm nb,min}<\varepsilon_{\rm inj,max}$ and ${\rm DM}_{\rm ej+nb}<{\rm DM}_{\rm host}$ are the most constraining for the range of parameters and models we considered. For $t_{\rm age}\gtrsim 10\,{\rm yr}$ and USSN progenitor, the allowed ($P_{i}$, $B_{\rm dip}$) parameter space is: ($1-10\,{\rm ms}$, $10^{12-14}\,{\rm G}$) for R1 and R2, and ($1-20\,{\rm ms}$, $10^{12-14}\,{\rm G}$) for R3. The corresponding parameter space for a CCSN progenitor is: ($1-2.5\,{\rm ms}$, $5\times 10^{12}-2\times 10^{14}\,{\rm G}$) for R1, ($1-20\,{\rm ms}$, $10^{12}-2\times 10^{14}\,{\rm G}$) for R2, and ($1-3\,{\rm ms}$, $5\times 10^{12}-3\times 10^{14}\,{\rm G}$) for R3. For both progenitor types and $t_{\rm age}\gtrsim 10\,{\rm yr}$, we found that the parameter space satisfying $L_{\rm sd}>10^{42}\,{\rm erg/s}$ overlaps with the allowed ($P_{i}$, $B_{\rm dip}$) parameter space, which indicates that the NS rotational energy is likely to be the primary energy reservoir for these FRB sources detected with persistent radio counterparts. As the flux decline rate is expected to be proportional to the current NS spindown timescale, $t_{\rm age}$ can also be inferred from detailed follow-up observations of the PRS.
@@
The localisation of R1 and R2 in a dwarf host galaxy with high specific star-formation rate (Chatterjee et al., 2017; Tendulkar et al., 2017) suggested a connection between FRBs and SLSN progenitors (Perley et al., 2016), as also previously pointed out by Murase et al. (2016). Modeling the quasi-steady nebular emission is therefore important to constrain the FRB source properties as the radio emission can be significantly absorbed in the MWN and SN ejecta at early times $t_{\rm age}\lesssim{\rm few\,yrs}$. SSA in the magnetized nebula and free-free absorption in the SN ejecta are the primary processes that can suppress the low-energy synchrotron flux for SNe progenitors with large ejecta mass and/or explosion energy. To this end, we compute the radio SEDs and light curves for synchrotron emission from the magnetized nebula using the code developed by Murase et al. (2015). We included the inverse Compton and synchrotron emission processes to solve the kinetic equations (see equation 3.2) to account for electromagnetic cascades in electron-positron pairs with the photons available from the magnetized nebula.
@@
We first considered the rotation-powered model with microscopic parameters $\epsilon_{B}=1-\epsilon_{e}=0.01$, $\gamma_{b}=10^{5}$, $q_{1}=1.5$ and $q_{2}=2.5$. We found that the observed radio flux for both R1 and R2 are consistent with an NS of $t_{\rm age}\approx 20\,{\rm yr}$ in a USSN progenitor. However, the observed persistent emission for R3 can only be explained with the CCSN model for a young NS aged $t_{\rm age}\approx 10\,{\rm yr}$. Although the synchrotron flux at lower energies is highly suppressed due to SSA, this effect is most prominent for R3 even at energies $\nu\lesssim 5\,{\rm GHz}$ and therefore suggests a larger $M_{\rm ej}$ (see Figure 3). Contrarily, the larger radio flux at $\nu\approx 1\,{\rm GHz}$ along with a harder energy spectra for R1 and R2 is better explained by the USSN progenitor with a smaller $M_{\rm ej}$. An increase in $B_{\rm dip}$ may not necessarily lead to a brighter radio emission as the NS spindown timescale $t_{\rm sd}\propto B_{\rm dip}^{-2}P_{i}^{2}$ becomes smaller, thereby resulting in more energy loss due to adiabatic expansion even at early times. From our radio SED estimates, we inferred that the observed PRS flux for R1 and R2 can be explained by an NS with magnetic field $B_{\rm dip}=(3-5)\times 10^{12}\,{\rm G}$ and initial spin period $P_{i}=1.5-3\,{\rm ms}$, whereas $B_{\rm dip}=5.5\times 10^{13}\,{\rm G}$ and $P_{i}=10\,{\rm ms}$ in case of R3 as shown in Figure 4.
@@
We also tested the alternate scenario whereby energy injection into the nebula is dominated by magnetar flares, corresponding to microscopic parameters $\epsilon_{B}=1-\epsilon_{e}=0.1$, $\gamma_{b}=10^{3}$ and $q_{1}=q_{2}=2$. For this case, we fixed the NS parameters to $P_{i}=300\,{\rm ms}$, $B_{\rm int}=5\times 10^{16}\,{\rm G}$ and $B_{\rm dip}=5\times 10^{14-15}\,{\rm G}$ and deduced the typical $t_{\rm age}$ such that the MWN can power the persistent radio emission. For R1 and R2, we found that the USSN progenitor model best explains the low-energy radio SED for $t_{\rm age}\approx 25\,{\rm yr}$ and $\approx 40\,{\rm yr}$, respectively. Large suppression in the observed radio flux for R3 implied that the source is likely to be a young NS with $t_{\rm age}\approx 12.5\,{\rm yr}$ located within a CCSN progenitor with larger ejecta mass as shown in Figure 7.
@@
The observed flux density of a PRS can also exhibit variability across long timescales that are likely to be intrinsic to the source. In a recent study, Rhodes et al. (2023) reported a $\sim 30\%$ reduction in the flux density of R1’s PRS at 1.3 GHz over three years (from 2019 to 2022) with the MeerKAT telescope. Similarly, Zhang et al. (2023) reported a $\sim 20\%$ reduction in the radio flux of R2’s PRS from their 3 GHz Very Large Array (VLA) observations in 2020 and 2021. Such long-term variation in the PRS radio flux can be instrumental in constraining the properties of the near-source region. Based on the predictions for radio light curves and their comparison with the current data in Figure 5, we found that the USSN and CCSN progenitor models are effectively indistinguishable for both R1 and R2. However, recent observations (Bruni et al., 2023) for R3’s PRS favor the CCSN model, in agreement with our radio SED results from Figures 3 and 4.
@@
While the persistent radio emission arises from the relativistic plasma, the DM contribution comes primarily from the cold plasma. If large luminosities for PRSs could be correlated with large ${\rm DM}_{\rm host}$ for these sources, that would have tremendous diagnostic power. The inferred ${\rm DM}_{\rm host}+{\rm DM}_{\rm ns}$ for these three FRBs indicates substantial contribution from the near-source plasma and circumgalactic medium. Using ionization and radiation transport models for the young NS, we estimated the contributions from the magnetized nebula and the SN ejecta to show that the latter dominates for typical NS parameters. In Figure 6, we obtained a minimum NS age $t_{\rm age,min}\sim 1-3\,{\rm yr}$ for the three FRB sources considered such that the SN ejecta is not significantly dense to overproduce observed DM. Non-attenuation of radio signal due to free-free absorption in the ejecta and synchrotron self-absorption in the nebula provides a stronger constraint with $t_{\rm age,min}\approx 10\,(8)\,[6.5]\,{\rm yr}$ for R1 (R2) [R3]. It should be noted that the inferred near-source DM is subject to some uncertainty as the pair multiplicity in our models can be uncertain (see Murase et al. 2015, for a detailed discussion).

### v2

@@ 5.3 Rotation powered vs magnetic powered models
Figure 7: Effect of NS age ($t_{\rm age}$) on the radio SED of persistent emission from FRB 20121102A (20190520B) [20201124A] source is shown in the left (center) [right] panel. The fluxes for synchrotron emission are shown for magnetar-flare-powered energy injection model with $\epsilon_{B}=0.1$, $\gamma_{b}=10^{3}$, $q_{1}=q_{2}=2.0$. For each source, we fix $P_{i}=300\,{\rm ms}$ and $B_{\rm dip}\sim(0.01-0.1)\,B_{\rm int}\approx 5\times 10^{14}-5\times 10^{15}\,{\rm G}$ to vary the NS age within a range that includes the best-fit $t_{\rm age}$. In each panel, the SED data from radio observations is shown with filled circles and the results for USSN/CCSN progenitors are shown using solid/dashed curves.
@@
Figure 8: Radio light curves for persistent emission associated with FRB 20121102A (20190520B) [20201124A] are shown in the left (center) [right] panel. Corresponding data at various frequencies is shown using unfilled circles for each source. The solid and dashed curves here denote the results for USSN and CCSN progenitors, respectively. We fix the NS parameters ($B_{\rm dip}$, $P_{i}$, $t_{\rm age}$) to their best-fit values as obtained from Figure 7 for the magnetar-flare-powered model.
@@
  FRB 20121102A USSN ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) $25\,{\rm yr}$ $3\times 10^{15}\,{\rm G}$
@@
FRB 20190520B USSN ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) $40\,{\rm yr}$ $5\times 10^{15}\,{\rm G}$
@@
FRB 20201124A USSN ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) $12.5\,{\rm yr}$ $5\times 10^{14}\,{\rm G}$
@@
Energy injected by the rotating magnetar drives the MWN and SN ejecta to evolve together, and the observed quasi-steady emission is powered by the synchrotron radiation generated from the magnetized nebula. For young rapidly rotating NS, the rotational energy is the primary energy reservoir that powers the magnetised nebula and it has been studied in detail for Galactic PWNe (Tanaka and Takahara, 2010, 2013). For a typical NS with mass $M_{\rm ns}=1.4\,M_{\odot}$ and radius $R_{\rm ns}=12\,{\rm km}$, the rotational energy is given by $\mathcal{E}_{\rm rot,i}\approx(1.9\times 10^{52}\,{\rm erg})\,P_{i,-3}^{-2}$. However, for a magnetar of age $t_{\rm age}\gg t_{\rm sd}\approx(0.12\,{\rm yr})\,B_{\rm dip,13}^{-2}P_{\rm i,-3}^{2}$, the central NS may spin down considerably over time such that its interior magnetic energy, $\mathcal{E}_{\rm B,int}\approx B_{\rm int}^{2}R_{\rm ns}^{3}/6=(3\times 10^{49}\,{\rm erg})\,B_{\rm int,16}^{2}$, is comparable or even larger than $\mathcal{E}_{\rm rot,i}$.
@@
In our analysis, we incorporate models for rotational energy injection proposed by Murase et al. (2016) and energy injection due to magnetar flares based on Margalit and Metzger (2018), to compute the total energy injected into the magnetar wind nebula (see equation 2.2.1). However, the microphysical parameters $\epsilon_{B}$, $\gamma_{b}$, $q_{1}$ and $q_{2}$ that determine the actual energy injection rate into the radiating electrons and positrons are given by equation (4). In the previous sections, we considered a rotation-powered model with $\epsilon_{B}=0.01$, $\gamma_{b}=10^{5}$, $q_{1}=1.5$ and $q_{2}=2.5$. The corresponding best-fit NS parameters ($t_{\rm age}$, $B_{\rm dip}$, $P_{i}$) are listed in Table 1 for USSN and CCSN progenitor models, based on the radio SEDs obtained for all three sources in Figures 3 and 4.
@@
We now consider the scenario where the energy injection from magnetar flares exceeds that from NS rotation ($\mathcal{E}_{\rm B,int}\gtrsim\mathcal{E}_{\rm rot,i}$), with $P_{i}=300\,{\rm ms}$, $B_{\rm int}=5\times 10^{16}\,{\rm G}$ and $B_{\rm dip}\sim 5\times 10^{14}-5\times 10^{15}\,{\rm G}$. We study the effect of varying $t_{\rm age}$ on the radio SED of synchrotron emission for the magnetar-flare-powered model with $\epsilon_{B}=0.1$, $\gamma_{b}=10^{3}$ and $q_{1}=q_{2}=2.0$. The results for FRB 20121102A/20190520B/20201124A are shown in the left/center/right panel of Figure 7, with solid/dashed curves denoting USSN/CCSN progenitors and filled circles corresponding to the radio data for each source. As for the rotation-powered model in Figure 3, the late-time ($t_{\rm age}\gg t_{\rm sd}$) and low-energy ($\nu\lesssim 1\,{\rm GHz}$) synchrotron flux is suppressed due to significant adiabatic losses and SSA, respectively. However, the high-energy radio flux typically attains smaller values compared to the rotation-powered model to then gradually fall off. This is expected as the energy injection into particles occurs at smaller $\gamma_{b}\sim 5\times 10^{3}$ for the magnetar-flare-powered model. We find that the USSN progenitor model better explains the low-energy ($\nu\lesssim 1\,{\rm GHz}$) radio SED for both FRBs 20121102A and 20190520B, corresponding to best-fit $t_{\rm age}\approx 25\,{\rm yr}$ and $\approx 40\,{\rm yr}$, respectively. On the contrary, the radio SED for FRB 20201124A is only explained by the CCSN model with a best-fit $t_{\rm age}\approx 12.5\,{\rm yr}$ while accounting for the large suppression in radio flux at smaller energies.
@@
For the magnetar-flare-powered model, we show the radio light curves for the persistent emission associated with each FRB source in Figure 8. We use the best-fit $B_{\rm dip}$ and $P_{i}$ values obtained from radio SED fits in Figure 7 for each FRB source, with the corresponding NS age set to $t_{\rm age}=25\,(40)\,[12.5]\,{\rm yr}$. In contrast to the results obtained in Figure 5 for the rotation-powered model, we find that the flux predictions from the USSN model are favored by the radio data, especially at smaller frequencies, for both FRBs 20121102A and 20190520B. In case of FRB 20201124A, the radio observations at smaller frequency $\nu\approx 6\,{\rm GHz}$ favor the CCSN model over the USSN progenitor model. These results are consistent with those obtained from the radio SED fits for the magnetar-flare-powered model ($\epsilon_{B}=0.1$, $\gamma_{b}=10^{3}$, $q_{1}=q_{2}=2.0$) in Figure 7. The best-fit NS parameters ($t_{\rm age}$, $B_{\rm dip}$) for the magnetar-flare-powered scenario with $P_{i}\approx 300\,{\rm ms}$ are listed in Table 2, for both USSN and CCSN models, based on the radio SED and light curve results shown in Figure 7 and 8, respectively.
6 Discussion
@@
Magnetars and rapidly rotating pulsars have been studied extensively as central engines of SLSNe and FRBs, as they can naturally explain the quasi-steady synchrotron emission from their nascent pulsar/magnetar wind nebulae. These central engines can provide a unified picture of SLSNe, stripped-envelope SNe and FRBs as well as long GRBs. Non-thermal nebular emission has been observed for many Galactic PWNe, which suggests that a significant fraction of the wind magnetic energy can be utilized for accelerating particles. Murase et al. (2016) proposed quasi-steady synchrotron emission as a potential probe of FRB progenitors and also explored their possible connection to pulsar/magnetar-driven SNe including SLSNe. Till date, five repeating FRBs namely 20121102A, 20190520B, 20201124A, 20240114A and 20190417A have been localized and associated with compact persistent radio sources (Chatterjee et al., 2017; Niu et al., 2022; Bruni et al., 2023, 2024; Moroianu et al., 2025).
@@
In this study, we consider the scenario whereby quasi-steady emission from such FRB sources is powered by young rapidly rotating magnetars that are embedded in a composite of magnetized wind nebula and SN ejecta. We have presented a detailed theoretical model here to numerically compute the radio SEDs for nebular emission by solving the time-dependent kinetic equations for photons and electrons and positrons in the MWN, including inverse Compton and synchrotron emission processes to account for electromagnetic cascades. To test SNe progenitor models, we considered two cases: (a) an ultra-stripped SNe with $M_{\rm ej}=0.1\,M_{\odot}$ and $E_{\rm sn}=10^{50}\,{\rm erg}$, and (b) a conventional core-collapse SNe with $M_{\rm ej}=3.0\,M_{\odot}$ and $E_{\rm sn}=10^{51}\,{\rm erg}$.
@@
While rotational energy from a young NS is likely to be the primary reservoir that powers magnetized nebula, energy released from the dissipation of interior magnetic fields (magnetar flares) can be significant at later times $t_{\rm age}\gg t_{\rm sd}$. We estimated the total energy injected into the MWN based on the NS rotational energy injection model proposed by Murase et al. (2016). For the magnetar flare energy injection model, we adopted parameters suggested by Beloborodov (2017). In addition to the SNe progenitor and energy injection models, synchrotron flux from the magnetized nebula is determined by NS parameters such as $B_{\rm dip}\sim 10^{12}-10^{15}\,{\rm G}$, $P_{i}\sim 1-30\,{\rm ms}$ and $t_{\rm age}\gtrsim t_{\rm obs}$.
@@
We first estimated the allowed combinations of these parameters using physical constraints from: (a) energy injected into the magnetized nebula ($\varepsilon_{\rm nb,min}<\varepsilon_{\rm inj,max}$), (b) near-source DM contribution from the SN ejecta and MWN (${\rm DM}_{\rm ej+nb}<{\rm DM}_{\rm host}$), (c) non-attenuation of radio signal due to free-free absorption and synchrotron self-absorption ($\tau_{\rm ff},\,\tau_{\rm sa}\lesssim 1$), and (d) size of the magnetized nebula ($R_{\rm nb}\lesssim 10\,{\rm pc}$). The corresponding results for $B_{\rm dip}$, $P_{i}$ and $t_{\rm age}$ are shown in Figure 2. We showed that $\varepsilon_{\rm nb,min}<\varepsilon_{\rm inj,max}$ and ${\rm DM}_{\rm ej+nb}<{\rm DM}_{\rm host}$ are the most constraining for the range of parameters and models considered. For $t_{\rm age}\gtrsim 10\,{\rm yr}$, the allowed ($P_{i}$, $B_{\rm dip}$) parameter space generally falls within ($1-10\,{\rm ms}$, $10^{12-14}\,{\rm G}$) for both USSN and CCSN progenitors across the three sources (FRBs 20121102A, 20190520B, 20201124A), with specific ranges detailed in the figure. For both progenitor types and $t_{\rm age}\gtrsim 10\,{\rm yr}$, we found that the parameter space satisfying $L_{\rm sd}>10^{42}\,{\rm erg/s}$ overlaps with the allowed ($P_{i}$, $B_{\rm dip}$) parameter space, which indicates that the NS rotational energy is likely to be the primary energy reservoir for these FRB sources detected with persistent radio counterparts. As the flux decline rate is expected to be proportional to the NS spindown timescale, $t_{\rm age}$ can also be inferred from detailed follow-up observations of the PRS.
@@
The localisation of FRBs 20121102A and 20190520B in dwarf host galaxies with high specific star-formation rates (Chatterjee et al., 2017; Tendulkar et al., 2017) suggested a connection between FRBs and SLSN progenitors (Perley et al., 2016), as also previously pointed out by Murase et al. (2016). Modeling the quasi-steady nebular emission is therefore important to constrain the FRB source properties, as the radio emission can be significantly absorbed in the MWN and SN ejecta at early times ($t_{\rm age}\lesssim{\rm few\,yrs}$). SSA in the magnetized nebula and free-free absorption in the SN ejecta are the primary processes that can suppress the low-energy synchrotron flux, particularly for SNe progenitors with large ejecta mass and/or explosion energy.
@@
Using our numerical code (based on Murase et al. 2015, 2021) that solves the kinetic equations including electromagnetic cascades, we computed radio SEDs and light curves. We first considered the rotation-powered model with microscopic parameters: $\epsilon_{B}=1-\epsilon_{e}=0.01$, $\gamma_{b}=10^{5}$, $q_{1}=1.5$ and $q_{2}=2.5$. We found that the observed radio flux for both FRBs 20121102A and 20190520B are consistent with an NS of $t_{\rm age}\approx 20\,{\rm yr}$ in a USSN progenitor. This progenitor provides a better match to the larger radio flux at $\nu\approx 1\,{\rm GHz}$ and the harder energy spectra observed for FRBs 20121102A and 20190520B. However, the observed persistent emission for FRB 20201124A, particularly the strong suppression at lower frequencies ($\nu\lesssim 5\,{\rm GHz}$ due to SSA), can only be explained with the CCSN model for a young NS aged $t_{\rm age}\approx 10\,{\rm yr}$, suggesting a larger $M_{\rm ej}$ (see Figure 3). From our radio SED estimates, we inferred NS parameters: for FRBs 20121102A and 20190520B, $B_{\rm dip}=(3-5)\times 10^{12}\,{\rm G}$ and $P_{i}=1.5-3\,{\rm ms}$; for FRB 20201124A, $B_{\rm dip}=5.5\times 10^{13}\,{\rm G}$ and $P_{i}=10\,{\rm ms}$ (as shown in Figure 4). An increase in $B_{\rm dip}$ does not necessarily lead to brighter radio emission due to shorter spindown timescales ($t_{\rm sd}\propto B_{\rm dip}^{-2}P_{i}^{2}$) causing greater adiabatic energy loss.
@@
We also tested the alternate scenario where energy injection is dominated by magnetar flares using $\epsilon_{B}=1-\epsilon_{e}=0.1$, $\gamma_{b}=10^{3}$, $q_{1}=q_{2}=2$, $P_{i}=300\,{\rm ms}$, $B_{\rm int}=5\times 10^{16}\,{\rm G}$ and $B_{\rm dip}=5\times 10^{14-15}\,{\rm G}$. From radio SEDs, we then inferred typical $t_{\rm age}$ such that the MWN can power the persistent radio emission. For FRBs 20121102A and 20190520B, the USSN progenitor model best explains the low-energy radio SED for $t_{\rm age}\approx 25\,{\rm yr}$ and $\approx 40\,{\rm yr}$, respectively. The large suppression in FRB 20201124A’s observed radio flux again implied a CCSN progenitor with larger ejecta mass, likely with $t_{\rm age}\approx 12.5\,{\rm yr}$ (see Figure 7).
@@
The observed flux density of a PRS can exhibit variability across long timescales that are intrinsic to the source. In a recent study, Rhodes et al. (2023) reported a $\sim 30\%$ reduction in the flux density of FRB 20121102A’s PRS at 1.3 GHz over three years (from 2019 to 2022) with the MeerKAT telescope. Similarly, Zhang et al. (2023) reported a $\sim 20\%$ reduction in the radio flux of FRB 20190520B’s PRS from their 3 GHz Very Large Array (VLA) observations in 2020 and 2021. Such long-term variation can be instrumental in constraining the properties of the near-source region. Based on our predicted radio light curves (Figure 5), the USSN and CCSN models are hard to distinguish for FRBs 20121102A and 20190520B based on current data. However, recent observations for FRB 20201124A’s PRS (Bruni et al., 2023) favor the CCSN model, consistent with our SED results.
@@
While the PRS arises from relativistic plasma, the near-source DM contribution comes primarily from cold plasma (dominated by SN ejecta in our models). The inferred ${\rm DM}_{\rm host}+{\rm DM}_{\rm ns}$ for these FRBs indicates substantial near-source contribution. Our models indicated a minimum NS age $t_{\rm age,min}\sim 1-3\,{\rm yr}$ for these sources such that the SN ejecta does not overproduce the observed DM (see Figure 6). However, non-attenuation of the radio signal due to free-free absorption in the ejecta and synchrotron self-absorption in the nebula provides a stronger constraint, yielding $t_{\rm age,min}\approx 10\,(8)\,[6.5]\,{\rm yr}$ for FRB 20121102A (20190520B) [20201124A]. The inferred near-source DM is subject to uncertainty from the pair multiplicity in our models (Murase et al., 2015).